The school you would like to attend costs $100,000. To help finance your education, you need to choose whether or not to sell any of your 500 shares of Apple stock you bought five years ago, 100 Apple bonds (each with a $1,000 face value and a 3.25% coupon rate) that are five years from their 10-year maturity date, or a combination of both. Provide the appropriate data and calculations that you would perform to make this decision.

Apple Stock price used: $105.11 (10/27/2014)
Answer #1

Option :

Coupon payment of the bonds = 100* 1000 * 3.25% = 3250

Total bond cash flow over 5 years = 3250 * 5 + 100,000 =116250

If you are selecting to keep stocks then you will have to forego 16250 ( 3250* 5 years) i.e. the coupon payment each year.

Explanation:

The shares were bought 5 years ago and will have appreciated by a great extent as on date. Since the share price at which it was bought is not given , hence it is difficult to calculate the annualized return for the stock. However, the returns of stock are tax free but the interest on bonds is taxable. Also, Apple being a company that has a high market capitalization, the stock will appreciate higher than the bond cashflows over the 5 years. Bonds offer fixed payments but stocks are a scope of multiplying the value. Hence, it is better to retain the stock and sell the bonds.